OpenStudy (anonymous):

reduce the equation into slope intercept form 6x-4y+11=0

OpenStudy (anonymous):

Slope intercept form is y=mx+b First you should move the constant and the mx over. This can be done by subtracting 6x from both sides and subtracting 11 from both sides. This will get you -4y = -6x - 11 Does that first step make sense?

OpenStudy (anonymous):

You will then need to get the y by itself. This can be done by diving -4 by both sides This will give you y = -6/4x - 11/4 = -3/2x - 11/4

OpenStudy (anonymous):

Sorry I got that wrong. The negatives will be positive = 3/2x + 11/3